# Walking the Tightrope of LLMs for Software Development: A Practitioners’ Perspective

### v1

@@ 3.1.1 Individual Level
Boosting code Development. A majority of participants (i.e., P1-P16, P18-P19, P22) mention LLMs speeding up software development by automating simple, repetitive, and tedious tasks, e.g.: 9 “You spend less time thinking about simple problems, and you spend more time thinking about complex problems” – P13 [Software Developer]. For instance, LLMs can provide the boilerplate files quickly, e.g.: 9 “I just started a new project. So okay, I’ll get the initial boilerplate done quickly. So I don’t need to go create multiple folders and the files and everything” – P10 [Software Engineer]. LLMs can also provide an extra boost to finish tasks quickly when stuck, e.g.: 9 “I try to show a code snippet and say [to ChatGPT]: ‘ I’ve developed it up to here, and I’d like to elaborate on it up to there. What would you do? Why?’” – P6 [Data Engineer]. Participants point out that, while developers could come up with those solutions, they would take more time, especially for novice developers, e.g.: 9 “As a software engineer, you can come up with all these solutions by yourself, but it would just take a longer period of time.” – P3 [Software Engineer]. In summary, this boost in developer productivity occurs when developers rely on LLMs’ capabilities, which save development time by reducing their effort and interruptions.
@@
LLMs Reducing Developers’ Effort. A majority of participants (i.e., P1-P18, P20-P22) described scenarios where LLMs work to lighten their efforts for simple, repetitive, or tedious tasks. This happens because software practitioners hand over the control to LLMs’ automation capabilities, e.g.: 9 “As I have this extension in Databricks, it makes it much easier. For example, [if] I miss a line, I miss a path, I miss a way of calling a class method, etc, it corrects it on the spot and automatically shows the button: ‘Do you want me to correct it and run the cell again?’.” – P6 [Data Engineer]. This mitigates interruptions in their development flow that would happen by doing it themselves, such as by filtering relevant content on Google, e.g.: 9 “If you go on to Google […] there’s a bunch of official documentation from different resources. It takes some cognitive load to pick one, and then it goes through.” – P11 [Web Developer].
@@
LLMs Saving Time. Most of the participants (i.e., P1-P18, P20, P22) also describe how LLMs contribute to time-saving for simple, repetitive, or tedious tasks. The perceived saving in developers’ time happens due to the reduced effort provided by LLMs when the developer hands over control to LLMs, e.g.: 9 “In the past, there were several sites, Stack Overflow […] From the moment you go looking for your question until you find it, you’ve already spent a lot of time.” – P6 [Data Engineer]. Thus, this time-saving is correlated to the reduced effort, requiring the developer to weaken the grip over development. Otherwise, developers might not notice significant time-saving, e.g., 9 “I only use it for syntax lookup. And I don’t do that very often […] and the time it really saves me is […] pretty much the exact same thing that I need to click into a website, load it and then scroll a little bit.” – P11 [Web Developer]. At the same time, participants do not feel confident about reducing the time required for task execution during planning, e.g., 9 “Something that will normally take me 3 days [to do without LLMs help], I’m still going to say 3 days. It’s not just because I have [Microsoft] Copilot doesn’t mean I’m going to say: ‘This will take me one day or 2 days’.” – P14 [Data Engineer].
@@
Gaining Learning Opportunities. Fourteen participants (i.e., P1-P4, P6-P7, P12, P14, P15, P17, P18, P20, P22) mentioned the potential of LLMs as an educational tool for software developers, as well as the learning opportunities from adopting LLMs. For instance, developers can search for the reason behind AI wrong suggestion, e.g.: 9 “I try to see if the wrong answer was because I didn’t give enough information […] I’ll see if the prompt didn’t make sense” – P1 [Software Developer]. LLMs can provide alternative solutions that can reinforce developers’ solution arsenal, e.g.: 9 “LLMs will give me an essentially different perspective. Maybe this is an alternative approach, a more concise way of doing things […] so then I can learn from them in this way.” – P17 [Data Analyst], and assist developers in programming language familiarisation, e.g.: 9 “I always used to search for Python functions in ChatGPT, […] Phind […] and also Perplexity.AI […] it was super helpful, especially to know the proper Python way of doing things, because each language has its own sort of style of writing code” – P2 [Software Engineer], and in getting domain knowledge, e.g.: 9 “GPT is definitely more helpful in […] adopting the background knowledge of a new area” – P11 [Web Developer].
@@
Benefit: [#1] Boosting code development, [#2] Reducing Effort, [#3] LLMs Saving Time, [#4] Gaining Learning Opportunities Reason:  LLMs Saving $\text{Time}^{\#1}$; LLMs Reducing $\text{Effort}^{\#1,\#3}$;  Mitigating $\text{Interruptions}^{\#1,\#2,\#3}$; Handing over the Control to $\text{LLMs}^{\#1,\#2,\#3}$; $\text{Automation}^{\#1,\#2,\#3}$; Reliance on $\text{LLMs}^{\#1,\#2,\#3,\#4}$; Balanced Control over $\text{LLMs}^{\#4}$; LLMs $\text{Capabilities}^{\#4}$
@@
Benefit: [#1] Boosting code development, [#2] Reducing Effort, [#3] LLMs Saving Time,[#4] Gaining Learning Opportunities,
@@
Maintaining Software Development Flow. Fifteen participants (i.e., P1, P4-P8, P12-P13, P15-P20, P22) presented different aspects that LLMs contribute to the software development flow. For instance, LLMs help developers to stay focused on the logic instead of typing when it autocompletes, e.g.: 9 “So it helps me with […] reducing the time I’m actually writing and giving me more space to think about the flow in general.” – P4 [Software Developer], fix syntax errors, e.g.: 9 “I really like to give ChatGPT the context […] As I have this extension in Databricks, it makes it much easier. For example, I miss a line, I miss a path, I miss a way of calling a class method, etc, it corrects it on the spot” – P6 [Data Engineer], or return relevant information, e.g.: 9 “You can just make a query like: ‘Hey, can you give me a query for this?’, [then] I can just copy [it and] paste it. [This basically] gives me more concentration for 2, 3 hours to solve that problem” – P12 [Software Engineer]. Participants also mention how browsing on Google may break the development flow, e.g.: 9 “When I didn’t have ChatGPT […] It was much more difficult to clear up my doubts by wandering around the internet. There’s a lot on the Internet […] several sites, Stack Overflow, for example. [But] for you to find your question and be able to find your specific answer took a while.” – P6 [Data Engineer]. When comparing AI pair programming and traditional pair programming, participants mentioned how traditional pair programming may come with interruption moments to searching for a solution, e.g.: 9 “Sometimes, when I have this question, and my colleague also does not know [how to get to] the answer, we have to look it up together in Stack Overflow [or something] like that.” – P8 [Software Engineer].
@@
Improving Developers’ Mental Model. A few participants (i.e., P6, P10, P13-P14) mention how LLMs can aid developers in improving their code understanding, e.g.: 9 “‘Why is it going on? Why is that method function [implemented in that way]?’” – P19 [Machine Learning Scientist]. When the developer responsible for a certain piece of code is busy, developers can ask questions to LLMs, e.g.:
9 “I have a friend who […] was the one who built the functions and all systems. Sometimes he’s unavailable [but] I [still] need someone to tell me what this [code] is about. When I read the code, [I cannot], it’s quite unclear [what the code is about]. [Then, my] 1st thought: ‘I need to check on documentation, and also some discussions on Stack Overflow’. [But then] I was thinking: ‘maybe I can use LLMs [to help understand the code]’. So, I copied and pasted the code [into Perplexity.AI] and I asked this LLM to explain that to me.” – P8 [Software Engineer].
Developers can also use LLMs to generate diagrams to improve their code understanding, e.g.: 9 “So if I have described an architecture. [I can say to the LLMs:] ‘Create the diagram for this architecture’. [Then] it creates the mermaid diagram, so it can easily be visualised how this architecture works.” – P10 [Software Engineer].
Safe Space. A few participants (i.e., P4, P7, P12, P15) mention how LLMs provide a safe space for sharing ideas and questions. For instance, software developers can ask as many questions as they need, e.g.: 9 “I feel like senior people don’t have a lot of time on hand for trivial issues […] [but for LLMs] there is no judgment [while asking questions], because you can ask dumb questions to ChatGPT or any [other] LLMs. And you can ask it as many times as you want, and you can get examples […] You can get customised responses.” – P15 [Researcher]. This seems to be motivated by the desire to avoid disturbing colleagues’ work, while becoming self-reliant, e.g.: 9 “Everyone’s very busy at work - the senior developers as well - you kind of - It’s not a set rule - but I feel I have a limited number of questions I can ask per day.” – P7 [BI Analyst]. In parallel, LLMs can offer certain levels of companionship experience, e.g.: 9 “We can interact with LLMs like a person. And it’s good to have that kind of companionship as well, like you can talk to it and say: ‘Hey, what do you think about this?’” – P12 [Software Engineer], and empathy, e.g.: 9 “There was one time when […] I imported some of the things Copilot sent me and tried them. When I went to Unity, it wouldn’t even run the project. […] [Copilot would reply]: ‘I’m sorry for your dissatisfaction, for your frustration, but let’s try to do it another way’.” – P4 [Software Developer].
@@ 3.1.2 Team Level
Mitigating Interruptions. Fifteen participants (i.e., P1, P4-P6, P10, P12-P21) described how LLMs lead to fewer interruptions. Participants showed a preference to LLMs over disturbing colleagues, e.g.: 9 “I’m always more comfortable using ChatGPT, Copilot. I do as little as possible to ask for help, so as not to disturb other people.” – P4 [Software Developer]. Participants also highlight the potential of LLMs supporting companies’ onboarding, which normally requires a lot of guidance from colleagues, e.g.: 9 “If you’re kind of new to the team, you would feel more inclined to kind of just break it down through Copilot or ChatGPT first. And then afterwards, if you still don’t understand it, then kind of go ask an experienced coworker” – P14 [Data Engineer].
Additional Assistance for Technical Questions. Many participants (i.e., P1-P4, P6, P8, P11-P13, P15, P17, P18, P22) mention about relying on LLMs for technical questions. LLMs are convenient in providing direction to software developers on how to approach tasks, e.g.: 9 “If I need some help […] I can get some idea about what I need to do regarding those things, if we need some best code practices, or something like that” – P18 [Software Developer]. LLMs can also suggest an alternative approach, acting as a second opinion, e.g.: 9 “Sometimes even the whole team is narrowed down into one specific way of doing stuff. So it’s a good thing there [to have a second opinion]” – P12 [Software Developer].
@@ 3.1.3 Organisation Level
Cost Saving due to LLMs. A few participants (i.e., P7, P9, P15-P16, P20-P22) highlight how LLMs can contribute to reducing costs. For instance, using LLMs for debugging can decrease the time to find the problem, e.g.: 9 “We are a small company, but we move at least one to 2 million dollars per week […] If there’s a production issue, then it’s like: ‘Oh, we need to solve this immediately’. And it [LLMs] saves a lot of time to find the bugs and to mitigate the problem. That [time spent searching for the problem] it costs, [and] you have put the money in somewhere else, where it should not be” – P12 [Software Engineer].
3.1.4 Society Level
Foster Entrepreneurship. A few participants (i.e., P10 and P21) mention the potential of LLMs encourage entrepreneurship. This occurs due to the great LLM prototyping capabilities, e.g.: 9 “Prototyping becomes much faster for these [startup] companies […] Even a non-coder […] can create just the rundown prototype, not a full, scalable solution. But just to showcase sort of a working front-end” – P21 [Researcher Developer], and boost in individual productivity, reducing demand for big teams, e.g.: 9 “It also makes it easier to create new companies […] because you can do more with fewer people. – P10 [Software Engineer]. At the same time, LLMs can provide basic information necessary to start a business, e.g.: 9 “It can help you with the basics of marketing, accounting, and law, so it can help you tap into the basics of all these things, so you don’t need to recruit a team first.” – P21 [Researcher Developer].
Consultant for Everyday Questions. A few participants (i.e., P6, P18, P21) highlight how LLMs emerge as a valuable assistant for everyday demands. This extends from simple questions, e.g.: 9 “In my day-to-day life, LLMs are more of a consultant for questions. For example, when I was arranging an interview with you, I thought: ‘My God, what’s the time zone difference between Brazil and Australia?’ LLM is a day-to-day consultant for basic questions.” – P6 [Data Engineer], and advanced tasks, e.g.: 9 “For general tasks, I have to plan a trip or something like that. I’m always like, sometimes like, go for ChatGPT and ask for [potential] schedules and everything. So, it’s easier for me [to use ChatGPT for searching relevant information].” – P18 [Software Developer].
RQ1. How do LLMs take software developers forward? We found benefits at the individual, team, organisation, and society levels. At the individual level, LLMs benefit developers by boosting software development, reducing effort, saving time, providing learning opportunities, maintaining software development flow, improving developers’ mental model, and providing a safe space. Most of the participants mentioned that LLMs boost software development, save time, and reduce effort.
3.2 RQ2. How do LLMs hold software developers back?
@@ 3.2.1 Individual Level
Slowing software development. Most of the participants (i.e., P1-P3, P5-P19, P21, P22) mention how using LLMs may slow software developers. For instance, they may need to start a new conversation when LLMs get stuck in the same wrong suggestion, e.g.: 9 “I usually started from scratch and gave him as much background and information as possible. Because when I continued the same conversation, it tended to continue the same mistake” – P6 [Data Engineer]. Due to the possibility of hallucinations happening, developers need to review suggestions, e.g., 9 “We don’t have that trust [that] everything [every output] is correct, because it’s not always correct, so we still have that gap like everything is not [always] correct.” – P18 [Software Developer].

### v2

@@ 3.1.1 Individual Level
Reducing developer effort with LLMs. A majority of participants (i.e., P1-P18, P20-P22) described scenarios where LLMs work to lighten their efforts for simple, repetitive, or tedious tasks. This happens because software practitioners hand over the control to LLMs’ automation capabilities, e.g.: “As I have this extension in Databricks, it makes it much easier. For example, [if] I miss a line, I miss a path, I miss a way of calling a class method, etc, it corrects it on the spot and automatically shows the button: ‘Do you want me to correct it and run the cell again?’.” – P6 [Data Engineer].
@@
Saving developer time with LLMs. Most of the participants (i.e., P1-P18, P20, P22) also describe how LLMs contribute to time-saving for simple, repetitive, or tedious tasks. The perceived saving in developers’ time happens due to the reduced effort provided by LLMs when the developer hands over control to LLMs, e.g.: “In the past, there were several sites, Stack Overflow […] From the moment you go looking for your question until you find it, you’ve already spent a lot of time.” – P6 [Data Engineer].
@@
Gaining learning opportunities. Fourteen participants (i.e., P1-P4, P6-P7, P12, P14, P15, P17, P18, P20, P22) mentioned the potential of LLMs as an educational tool for software developers, as well as the learning opportunities from adopting LLMs. For instance, developers can search for the reason behind AI’s wrong suggestion, e.g.: “I try to see if the wrong answer was because I didn’t give enough information […] I’ll see if the prompt didn’t make sense” – P1 [Software Developer]. LLMs can provide alternative solutions that can reinforce developers’ solution arsenal, e.g.: “LLMs will give me an essentially different perspective. Maybe this is an alternative approach, a more concise way of doing things” – P17 [Data Analyst].
@@
Benefit: [#1] Reducing effort, [#2] LLMs saving time, [#3] Gaining learning opportunities Reason:   LLMs Reducing $\text{Effort}^{\#2}$;  Mitigating $\text{interruptions}^{\#1,\#2}$; Handing over the Control to $\text{LLMs}^{\#1,\#2}$; $\text{Automation}^{\#1,\#2}$; Reliance on $\text{LLMs}^{\#1,\#2,\#3}$; Balanced control over $\text{LLMs}^{\#3}$; LLMs $\text{capabilities}^{\#3}$
@@
Benefit: [#1] Reducing effort, [#2] LLMs saving time,[#3] Gaining learning opportunities,
@@
Maintaining developer flow. Fifteen participants (i.e., P1, P4-P8, P12-P13, P15-P20, P22) presented different aspects that LLMs contribute to the software development flow. For instance, LLMs help developers to stay focused on the logic instead of typing when it autocompletes, e.g.: “So it helps me with […] reducing the time I’m actually writing and giving me more space to think about the flow in general.” – P4 [Software Developer].
@@
Improving developer mental models. A few participants (i.e., P6, P10, P13-P14) mention how LLMs can aid developers in improving their code understanding, e.g.: “‘Why is it going on? Why is that method function [implemented in that way]?’” – P19 [Machine Learning Scientist].
Safe space. A few participants (i.e., P4, P7, P12, P15) mention how LLMs provide a safe space for sharing ideas and questions. For instance, software developers can ask as many questions as they need, e.g.: “I feel like senior people don’t have a lot of time on hand for trivial issues […] [but for LLMs] there is no judgment [while asking questions], because you can ask dumb questions to ChatGPT or any [other] LLMs. And you can ask it as many times as you want, and you can get examples […] You can get customised responses.” – P15 [Researcher]. This seems to be motivated by the desire to avoid disturbing colleagues’ work, while becoming self-reliant, e.g.: “Everyone’s very busy at work - the senior developers as well - you kind of - It’s not a set rule - but I feel I have a limited number of questions I can ask per day.” – P7 [BI Analyst]. In parallel, LLMs can offer certain levels of companionship experience, e.g.: “We can interact with LLMs like a person. And it’s good to have that kind of companionship as well, like you can talk to it and say: ‘Hey, what do you think about this?’” – P12 [Software Engineer], and empathy, e.g.: “There was one time when […] I imported some of the things Copilot sent me and tried them. When I went to Unity, it wouldn’t even run the project. […] [Copilot would reply]: ‘I’m sorry for your dissatisfaction, for your frustration, but let’s try to do it another way’.” – P4 [Software Developer].
@@ 3.1.2 Team Level
Mitigating interruptions. Fifteen participants (i.e., P1, P4-P6, P10, P12-P21) described how LLMs lead to fewer interruptions. Participants showed a preference to LLMs over disturbing colleagues, e.g.: “I’m always more comfortable using ChatGPT, Copilot. I do as little as possible to ask for help, so as not to disturb other people.” – P4 [Software Developer]. Participants also highlight the potential of LLMs supporting companies’ onboarding, which normally requires a lot of guidance from colleagues, e.g.: “If you’re kind of new to the team, you would feel more inclined to kind of just break it down through Copilot or ChatGPT first. And then afterwards, if you still don’t understand it, then kind of go ask an experienced coworker” – P14 [Data Engineer].
@@ 3.1.3 Organisation Level
Cost savings due to LLMs. A few participants (i.e., P7, P9, P15-P16, P20-P22) highlight how LLMs can contribute to reducing costs. For instance, using LLMs for debugging can decrease the time to find the problem, e.g.: “We are a small company, but we move at least one to 2 million dollars per week […] If there’s a production issue, then it’s like: ‘Oh, we need to solve this immediately’. And it [LLMs] saves a lot of time to find the bugs and to mitigate the problem. That [time spent searching for the problem] it costs, [and] you have put the money in somewhere else, where it should not be” – P12 [Software Engineer].
3.1.4 Societal Level
Foster Entrepreneurship. A few participants (i.e., P10 and P21) mention the potential of LLMs to encourage entrepreneurship. This occurs due to the great LLM prototyping capabilities, e.g.: “Prototyping becomes much faster for these [startup] companies […] Even a non-coder […] can create just the rundown prototype, not a full, scalable solution. But just to showcase sort of a working front-end” – P21 [Researcher Developer], and boost in individual productivity, reducing demand for big teams, e.g.: “It also makes it easier to create new companies […] because you can do more with fewer people. – P10 [Software Engineer]. At the same time, LLMs can provide basic information necessary to start a business, e.g.: “It can help you with the basics of marketing, accounting, and law, so it can help you tap into the basics of all these things, so you don’t need to recruit a team first.” – P21 [Researcher Developer].
RQ1. How do LLMs help software developers move forward? Our findings reveal benefits at the individual (nn=6), team (nn=1), organizational (nn=1), and societal (nn=1) levels. Specifically, LLMs assist software developers to sustaining cognitive flow, enhancing mental models, and offering a psychologically safe and continuous learning environment. LLMs can also reduce developer effort and save developer time.
3.2 RQ2. How do LLMs create an imbalance in software development?
@@ 3.2.1 Individual Level
Frictions in LLM-assisted software development. Most of the participants (i.e., P1-P3, P5-P19, P21, P22) mention how using LLMs may slow software developers. For instance, they may need to start a new conversation when LLMs get stuck in the same wrong suggestion, e.g.: “I usually started from scratch and gave him as much background and information as possible. Because when I continued the same conversation, it tended to continue the same mistake” – P6 [Data Engineer].